# JUHAN ARU, PHILÉMON BORDEREAU

### v1

@@ Appendix A Existence of level-lines for compactified GFF on crossing admissible domains
Let $\Phi_{\mathbb{A}_{p}}^{0}$ be a (zero Dirichlet) GFF in $\mathbb{A}_{p}$ and $\Phi_{\mathbb{A}_{p}}^{g_{b_{\text{in}}}}:=\Phi_{\mathbb{A}_{p}}^{0}+g_{b_{\text{in}}}$. Then there exists a local set coupling $(\Phi_{\mathbb{A}_{p}}^{g_{b_{\text{in}}}},\eta,h_{\eta})$ where $h_{\eta}$ is multivalued in $\mathbb{A}_{p}\setminus\eta$ with constant branches equal to $\lambda+2n\lambda$, and $\eta$ is an annulus $\mathrm{SLE}_{4}$ in $\mathbb{A}_{p}$, i.e. a Loewner chain for the annulus Loewner equation [Zha04] with driving function corresponding to a semimartingale with an explicit drift. The winding of $\eta$ at $y$ is almost surely equal to $b_{\text{in}}+\alpha$, with $\alpha$ the representative of $(\arg y-\arg x)/(2\pi)$ in $[0,1)$.
The proof is very similar to the case of the usual level line coupling in [SS13, Dub09] and relies on finding the right martingales to then use Itô’s formula: a sufficient condition for the law of the GFF to be equal conditionally on $\eta_{t}$ to a GFF in $\mathbb{A}_{p}\setminus\eta_{t}$ with mean given by $g_{b_{\text{in}},t}:=f_{t}^{-1}(g_{b_{\text{in}}})$, where $f_{t}:\mathbb{A}_{p}\to\mathbb{A}_{p-t}$ is the mapping-out function (recall that the modulus plays the role of the time-parametrization), is that for all $z\in\mathbb{A}_{p}$, $\left\{g_{t}(z)\right\}_{0\leq t\leq p}$ is a (continuous) uniformly bounded martingale with quadratic cross variation given by $G_{\mathbb{A}_{p}}(z,z^{\prime})-G_{\mathbb{A}_{p-\cdot}}(z,z^{\prime})$. In other words, $G_{\mathbb{A}_{p-\cdot}}(z,z^{\prime})$ should be a process of finite variation with $g_{\cdot}(z)g_{\cdot}(z^{\prime})-G_{\mathbb{A}_{p-\cdot}}(z,z^{\prime})$ a martingale. This guarantees that the characteristic function of the two GFFs mentioned above agree. Itô’s formula then provides differential equations that the drift of the driving function should satisfy in order for the conditions above to be met.
In fact, the form of the coupling that is established in [IK13, Proposition 10] is weaker than a local set coupling, but this is because their set-up (see their Theorem 2) encompasses general flow-lines where $h_{\eta}$ is possibly unbounded in a neighborhood of $\eta$, which is not the case above; in the case of level lines the local set coupling can be achieved as in [SS13].
When $D$ is $n$-connected with $n\geq 1$, existence of level-lines follows by absolute continuity: the argument is almost identical to the proof of Lemma 15 in [ASW19]. By conformal invariance we can take $D$ as an annulus $\mathbb{A}_{p}$ minus a set $K$ given by a finite union of closed disks. Let also $g_{D}$ be a multivalued harmonic function in $D$ with locally constant boundary conditions across $x$ and $y$, a change of fixed branch of $2\lambda b_{\text{in}}$ from the outer to the inner boundary of $D$, arbitrary values on $\partial K$, and such that $g(re^{i\theta+i2\pi})=g(re^{i\theta})+2\lambda$. Given $D^{\prime}\subset D$ a test domain, one can define the level-line $\eta$ of $\Phi_{D}^{g_{D}}:=\Phi_{D}^{0}+g_{D}$ up until its exit time of $D^{\prime}$ as the level-line in $D^{\prime}$ of a GFF $\Phi_{D^{\prime}}^{0}+g_{D}|_{D^{\prime}}$. Uniqueness of this level-line implies that they are compatible so that $\eta$ is well-defined (and actually a measurable function of the field). Similarly, if all branches of $g_{D}$ are such that |$\absolutevalue{g_{D}}\geq\lambda$ on $\partial K$, then almost surely $\eta$ doesn’t hit $\partial K$: the argument is here almost identical to Lemma 16 in [ASW19]. It therefore almost surely ends at $y$ with a winding described by $b_{\text{in}}$ since for all $D^{\prime}\subset D$ a test domain, the law of $\eta$ up to its exit time of $D^{\prime}$ is absolutely continuous with respect to that of the level-line $\tilde{\eta}$ of $\Phi_{\mathbb{A}_{p}}^{g_{b_{\text{in}}}}$, which ends at $y$ with winding described by $b_{\text{in}}$ on the event of positive probability that it does not exit $D^{\prime}$.
Appendix B Proof of Lemma 10
Proof of Lemma 10.
@@ Appendix B Proof of Lemma 10
But on the other hand, denoting $R(x)=\mathrm{d}\mathbb{P}_{X}/\mathbb{P}_{Y_{1}}(x)$, we have that

### v2

@@ Appendix A Existence of level-lines for compactified GFF on crossing admissible domains
Let $(D,x,y)$ be a crossing admissible domain, $\mathbf{b}\in\Lambda^{n+1}$ and $\Phi_{D}^{g_{\mathbf{b}}}=\Phi_{D}^{0}+g_{\mathbf{b}}$. There exists a local set coupling $(\Phi_{D}^{g_{\mathbf{b}}},\eta,h_{\eta})$ where $h_{\eta}$ is multivalued in $D\setminus\eta$ with constant branches equal to $\lambda+2n\lambda$ on the event $\mathcal{A}_{\mathbf{b}}$. $\eta$ is almost surely a simple curve from $x$ to $y$ in $\mathbb{A}_{p}$, measurable w.r.t. $\Phi_{D}^{g_{\mathbf{b}}}$, and its winding at $y$ is almost surely equal to $b_{\text{in}}+\alpha$, with $\alpha$ the representative of $(\arg y-\arg x)/(2\pi)$ in $[0,1)$.
Proof.
By conformal invariance we can assume that $D$ is given by an annulus $\mathbb{A}_{p}$ minus a set $K$ corresponding to a finite union of closed disks whose boundaries are given by $\mathcal{B}_{j},1\leq j\leq n,n\geq 0$. Given $D^{\prime}\subset D$ a test domain, one can define the level-line $\eta$ of $\Phi_{D}^{g_{D}}\coloneqq\Phi_{D}^{0}+g_{D}$ up until its exit time of $D^{\prime}$ as the level-line in $D^{\prime}$ of a GFF $\Phi_{D^{\prime}}^{g_{D}|_{D^{\prime}}}\coloneqq\Phi_{D^{\prime}}^{0}+g_{D}|_{D^{\prime}}$. Uniqueness and measurability of this level-line w.r.t the latter GFF, as proved in the seminal works [12, 35], imply that this construction is compatible between different test domains so that $\eta$ is well-defined and actually a measurable function of $\Phi_{D}^{g_{D}}$. Similarly, since all branches of $g_{D}$ are such that $g_{D}$ is constant on each connected component of $\partial D\setminus\left\{x,y\right\}$ and equal to $(2k+1)\lambda$ for some $k\in\mathbb{Z}$, then almost surely $\eta$ doesn’t hit $\partial D\setminus\left\{y\right\}$: for any simply connected domain $D^{\prime}\subset D$, any point $y^{\prime}$ of $(\partial D\cap\partial D^{\prime})\setminus\left\{x,y\right\}$ admits a neighbourhood $J$ in $D^{\prime}$ such that $g_{\mathbf{b}}|_{J\cap\partial D^{\prime}}$ is equal to $(2k+1)\lambda$ for some $k\in\mathbb{Z}$, so by [32, Lemma 4.2] the probability that the level-line of $\Phi_{D^{\prime}}^{g_{D}|_{D^{\prime}}}$ converges to a point in $J\cap\partial D^{\prime}$ before exiting $D^{\prime}$ is zero. On the other hand, if $\partial D^{\prime}\ni y$, then there is a neighbourhood $J$ of $y$ in $D^{\prime}$ such that $g_{D}|_{J\cap\partial D^{\prime}}$ is piecewise constant equal to $(2k-1)\lambda$ (resp. $(2k+1)\lambda$) on the part of $J\cap\partial D^{\prime}$ in the clockwise direction (resp. anticlockwise) from $y$, for some $k\in\mathbb{Z}$. This $k$ is equal to zero if and only if the winding of $D^{\prime}$ from $x$ to $y$ is given by [^13] In the sense that any curve $\gamma$ from $x$ to $y$ in $D^{\prime}$ has winding given by $b_{\text{in}}+\alpha$. $b_{\text{in}}+\alpha$, so if it is not the case by [32, Lemma 4.2] again the probability that the level-line of $\Phi_{D^{\prime}}^{g_{D}|_{D^{\prime}}}$ converges to $y$ before exiting $D^{\prime}$ is zero. We conclude by the construction of the level-line in $D$ and the union bound (for instance, considering all $D^{\prime}$ corresponding the interior of finite unions of dyadic squares such that $D^{\prime}$ is simply connected and contains $x$) that $\eta$ almost surely ends at $y$ with winding given by $b_{\text{in}}+\alpha$. ∎
In [19], this result is established in the specific case $D=\mathbb{A}_{p}$, and it is further shown that $\eta$ is an annulus $\mathrm{SLE}_{4}$ in $\mathbb{A}_{p}$, i.e. a Loewner chain for the annulus Loewner equation [46] with driving function corresponding to a semimartingale with an explicit drift.
Appendix B Proof of Proposition 1.2
We remind the proposition:
Proposition 1.2.
For $D$ a simply connected proper domain of $\mathbb{C}$ and $\mathcal{B}_{1},\mathcal{B}_{2}$ disjoint simple crosscuts such that $\mathrm{dist}(\mathcal{B}_{1},\mathcal{B}_{2})>0$ and $\mathcal{B}_{1}$ is smooth, the following identity holds:
$\exp(-m_{D}(\mathcal{B}_{1},\mathcal{B}_{2}))=\mathrm{det}_{F}(I_{\mathcal{B}_{1}}-(G_{D\setminus\mathcal{B}_{2}})_{\mathcal{B}_{1}}^{\mathcal{B}_{1}}(H_{\partial(D\setminus\mathcal{B}_{1})}-H_{\partial(D\setminus(\mathcal{B}_{1}\cup\mathcal{B}_{2}))})_{\mathcal{B}_{1}}^{\mathcal{B}_{1}}).$
Proof.
The identity holds true when $\mathcal{B}_{2}$ is smooth and we have to extend it when $\mathcal{B}_{2}$ is merely continuous. Since smooth crosscuts are dense in the space of continuous crosscuts with the Hausdorff topology, it is enough to argue continuity in $\mathcal{B}_{2}$ of both sides of the identity. More precisely, fix two distinct endpoints $x,y$ on $\partial D$ and consider $\mathcal{X}$ the set of smooth crosscuts $\mathcal{B}_{2}$ having these two endpoints and disjoint from $\mathcal{B}_{1}$, equipped with the Hausdorff metric. On the one hand, the map $\mathcal{B}_{2}\mapsto m_{D}(\mathcal{B}_{1},\mathcal{B}_{2})$ defined on $\mathcal{X}$ is continuous (for instance by dominated convergence). On the other hand, since $A\mapsto\mathrm{det}_{\mathrm{F}}(I-A)$ is continuous over trace-class operators (with trace-class topology, [40, Theorem 3.4]) the result follows provided we can prove the following:
Claim.
The map $\mathcal{B}_{2}\mapsto(G_{D\setminus\mathcal{B}_{2}})_{\mathcal{B}_{1}}^{\mathcal{B}_{1}}(H_{\partial(D\setminus\mathcal{B}_{1})}-H_{\partial(D\setminus(\mathcal{B}_{1}\cup\mathcal{B}_{2}))})_{\mathcal{B}_{1}}^{\mathcal{B}_{1}}$ is continuous from $\mathcal{X}$ to the set of trace-class operators on $L^{2}(\mathcal{B}_{1})$.
Indeed, consider a sequence of crosscuts $(\eta_{k})_{k\geq 1}\subset\mathcal{X}$ converging to some $\eta_{\infty}\in\mathcal{X}$. Denote
$A_{k}:=(G_{D\setminus\eta_{k}})_{\mathcal{B}_{1}}^{\mathcal{B}_{1}}(H_{\partial(D\setminus\mathcal{B}_{1})}-H_{\partial(D\setminus(\mathcal{B}_{1}\cup\eta_{k}))})_{\mathcal{B}_{1}}^{\mathcal{B}_{1}}\quad\quad\text{ for }k\in\mathbb{N}\cup\left\{\infty\right\}.$
These operators have kernel
[…]
Appendix C Proof of Lemma 2.8
We remind the lemma:
Lemma 2.8.
Let $X,Y_{1},Y_{2}$ be random variables on the same probability space, taking values in a Polish space $E$, with associated distributions $\mathbb{P}_{X},\mathbb{P}_{Y_{1}},\mathbb{P}_{Y_{2}}$ respectively. Denote also by $\mathbb{P}_{Y_{1}|Y_{2}=y_{2}}$ the conditional law of $Y_{1}$ on $Y_{2}=y_{2}$} (similarly for $\mathbb{P}_{Y_{2}|Y_{1}=x}$) and assume that $\mathbb{P}_{X}\ll\mathbb{P}_{Y_{1}|Y_{2}=y_{2}}$, for $\mathbb{P}_{Y_{2}}$-almost every $y_{2}\in E$. Then the following equality between Radon-Nikodym derivatives holds $\mathbb{P}_{X}$-a.s.
Proof.
@@
But on the other hand, denoting $R(x)=\mathrm{d}\mathbb{P}_{X}/\mathrm{d}\mathbb{P}_{Y_{1}}(x)$, we have that